OpenStudy (anonymous):

Solve for x. 3^(2x)+3^(x+1)-4=0 Please help!

OpenStudy (dumbcow):

rewrite it as a quadratic \[(3^x)^2 + 3(3^x) -4 = 0\] u = 3^x \[u^2 +3u -4 = 0\] solve for u first , then plug 3^x back in to solve for x